# FiniteTimeBibliography.bib

### v1

@@ D.3 Quantum energy-band thermalization for accessible time scales
From Eq. (64), using the decomposition of eigenspaces into the two energy shells $\Sigma_{d_{s}}$ and $\Sigma_{\delta d}$ as in Eq. (188), with respective projectors $\hat{\Pi}_{s}(E)$ and $\hat{\Pi}_{\delta}(E)$, we get
$\frac{1}{d}\sum_{\mathcal{H}_{s}(E)\subseteq\Sigma_{d_{s}}}\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}_{s}(E)\right\}^{2}\right]+\frac{1}{d}\sum_{\mathcal{H}_{\delta}(E)\subseteq\Sigma_{\delta d}}\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}_{\delta}(E)\right\}^{2}\right]<\epsilon^{2}.$ (212)
@@
$\frac{1}{d}\sum_{\mathcal{H}_{\delta}(E)\subseteq\Sigma_{\delta d}}\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}_{\delta}(E)\right\}^{2}\right]<\epsilon^{2}.$ (213)
@@
$\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}_{s}(E)\right\}^{2}\right]<\lambda^{2},$ (214)
@@
$\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}_{\delta}(E)\right\}^{2}\right]\geq\lambda^{2}.$ (215)
Eq. (215) further implies
$\frac{1}{d}\sum_{\mathcal{H}_{\delta}(E)\subseteq\Sigma_{d}}\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}_{\delta}(E)\right\}^{2}\right]\geq\frac{\lambda^{2}}{d}\left(\sum_{\mathcal{H}_{\delta}(E)\subseteq\Sigma_{\delta d}}1\right).$ (216)
Combining this with Eq. (213), we get
$\left(\sum_{\mathcal{H}_{\delta}(E)\subseteq\Sigma_{\delta d}}1\right)\leq\frac{\epsilon^{2}}{\lambda^{2}}d,$ (217)
which is Eq. (189).
@@
$\left\lvert\int\mathrm{d}t\ w(t)\operatorname{Tr}[\hat{\rho}(t)\hat{\Pi}_{A}]-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\right\rvert=\left\lvert\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}\rangle,\lvert E_{m}\rangle\in\Sigma_{d}\end{subarray}}\widetilde{w}(E_{n}-E_{m})\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\langle E_{m}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\lvert E_{n}\rangle\right\rvert.$ (218)
To simplify our notation, we will take $\lvert E_{n}\rangle,\lvert E_{m}\rangle\in\Sigma_{d}$ as given in what follows, without explicitly writing out this condition in the sums. Separating Eq. (218) into contributions from within and outside the bandwidth $\Delta E$ (noting that $\hat{\mathds{1}}$ has no off-diagonal matrix elements) and using the triangle inequality, we get
$\left\lvert\int\mathrm{d}t\ w(t)\operatorname{Tr}[\hat{\rho}(t)\hat{\Pi}_{A}]-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\right\rvert\leq$ $\left\lvert\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\widetilde{w}(E_{n}-E_{m})\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\langle E_{m}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\lvert E_{n}\rangle\right\rvert$
$+\left\lvert\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\widetilde{w}(E_{n}-E_{m})\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\langle E_{m}\rvert\hat{\Pi}_{A}\lvert E_{n}\rangle\right\rvert.$ (219)
@@
$\leq\sqrt{\left(\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\left\lvert\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\right\rvert^{2}\right)\operatorname{Tr}\left\{\left[\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right]^{2}_{(\Sigma_{d},\Delta E)}\right\}}$ (220)
@@
$\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\left\lvert\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\right\rvert^{2}\leq\operatorname{Tr}[\hat{\rho}^{2}].$ (221)
@@
$\left\lvert\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\widetilde{w}(E_{n}-E_{m})\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\langle E_{m}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\lvert E_{n}\rangle\right\rvert<\epsilon\sqrt{d\operatorname{Tr}[\hat{\rho}^{2}]}.$ (222)
@@
$\leq w_{0}\sqrt{\left\{\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\left\lvert\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\right\rvert^{2}\right\}\left\{\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\left\lvert\langle E_{m}\rvert\hat{\Pi}_{A}\lvert E_{n}\rangle\right\rvert^{2}\right\}}$ (223)
@@
$\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\left\lvert\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\right\rvert^{2}$ $\leq\operatorname{Tr}[\hat{\rho}^{2}]$ (224)
$\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\left\lvert\langle E_{m}\rvert\hat{\Pi}_{A}\lvert E_{n}\rangle\right\rvert^{2}$ $\leq\operatorname{Tr}\left[\left\{\hat{\Pi}_{A}\hat{\Pi}_{\Sigma_{d}}\right\}^{2}\right].$ (225)
@@
$\operatorname{Tr}\left[\left\{\hat{\Pi}_{A}\hat{\Pi}_{\Sigma_{d}}\right\}^{2}\right]\leq\operatorname{Tr}[\hat{\Pi}_{A}\hat{\Pi}_{\Sigma_{d}}]=d\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}},$ (226)
@@
$\left\lvert\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\widetilde{w}(E_{n}-E_{m})\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\langle E_{m}\rvert\hat{\Pi}_{A}\lvert E_{n}\rangle\right\rvert\leq w_{0}\sqrt{d\operatorname{Tr}[\hat{\rho}^{2}]\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}}.$ (227)
Inserting Eqs. (222) and (227) into Eq. (219), we get Eq. (66).
@@
Noting that $\lambda_{k}[w]\in\mathbb{R}$ [defined in Eq. (74)], it is convenient to separate the basis states in $\mathcal{B}$ into a set of states $\mathcal{B}_{+}$ in which $\lambda_{k}[w]\geq 0$, numbering $n_{+}$, and a set $\mathcal{B}_{-}$ in which $\lambda_{k}[w]<0$, with $n_{-}$ states:
$\mathcal{B}_{+}$ =$=\{\lvert k\rangle\in\mathcal{B}:\ \lambda_{k}[w]\geq 0\},\;\;\ \lvert\mathcal{B}_{+}\rvert=n_{+},$ (228)
$\mathcal{B}_{-}$ =$=\{\lvert k\rangle\in\mathcal{B}:\ \lambda_{k}[w]<0\},\;\;\ \lvert\mathcal{B}_{+}\rvert=n_{-}.$ (229)
@@
$\hat{\rho}_{+}$ $\equiv\frac{1}{n_{+}}\sum_{\lvert k\rangle\in\mathcal{B}_{+}}\lvert k\rangle\langle k\rvert,$ (230)
$\hat{\rho}_{-}$ $\equiv\frac{1}{n_{-}}\sum_{\lvert k\rangle\in\mathcal{B}_{-}}\lvert k\rangle\langle k\rvert,$ (231)
@@
$\frac{1}{n_{\pm}}\sum_{k\in\mathcal{B}_{\pm}}\left\lvert\lambda_{k}[w]\right\rvert<\left(\epsilon+w_{0}\sqrt{\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}}\right)\sqrt{\frac{d}{n_{\pm}}},$ (232)
@@
$\frac{1}{d}\sum_{k\in\mathcal{B}_{\pm}}\left\lvert\lambda_{k}[w]\right\rvert<\left(\epsilon+w_{0}\sqrt{\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}}\right)\frac{\sqrt{n_{+}}+\sqrt{n_{-}}}{\sqrt{d}}\leq\left(\epsilon+w_{0}\sqrt{\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}}\right)\sqrt{2},$ (233)
@@
$f\lambda\leq\frac{1}{d}\sum_{k\in\mathcal{B}_{\pm}}\left\lvert\lambda_{k}[w]\right\rvert,$ (234)
which, on substitution into Eq. (233), gives Eq. (76).
@@ D.4 Global thermalization from a single initial state
$+\frac{1}{\operatorname{Tr}[\hat{\Pi}_{A}]}\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\widetilde{w}_{+}(E_{n}-E_{m})\left\lvert\langle E_{n}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle\hat{\mathds{1}}\right)\lvert E_{m}\rangle\right\rvert^{2}$ (235)
@@
$\int\mathrm{d}t\ w_{+}(t)\operatorname{Tr}[\hat{\rho}_{A}(t)\hat{\Pi}_{A}]-\langle\hat{\Pi}_{A}\rangle\geq\frac{1}{\operatorname{Tr}[\hat{\Pi}_{A}]}\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\widetilde{w}_{+}(E_{n}-E_{m})\left\lvert\langle E_{n}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle\hat{\mathds{1}}\right)\lvert E_{m}\rangle\right\rvert^{2}$ (236)
@@
$\frac{1}{D}\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\left\lvert\langle E_{n}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle\hat{\mathds{1}}\right)\lvert E_{m}\rangle\right\rvert^{2}<\frac{\operatorname{Tr}[\hat{\Pi}_{A}]}{WD}\left[\int\mathrm{d}t\ w_{+}(t)\operatorname{Tr}[\hat{\rho}_{A}(t)\hat{\Pi}_{A}]-\langle\hat{\Pi}_{A}\rangle\right].$ (237)

### v2

@@ D.3 Quantum energy-band thermalization for accessible time scales
From Eq. (64), using the decomposition of eigenspaces into the two energy shells $\Sigma_{d_{s}}$ and $\Sigma_{\delta d}$ as in Eq. (C.1), with respective projectors $\hat{\Pi}_{s}(E)$ and $\hat{\Pi}_{\delta}(E)$, we get
$\frac{1}{d}\sum_{\mathcal{H}_{s}(E)\subseteq\Sigma_{d_{s}}}\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}_{s}(E)\right\}^{2}\right]+\frac{1}{d}\sum_{\mathcal{H}_{\delta}(E)\subseteq\Sigma_{\delta d}}\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}_{\delta}(E)\right\}^{2}\right]<\epsilon^{2}.$ (D.17)
@@
$\frac{1}{d}\sum_{\mathcal{H}_{\delta}(E)\subseteq\Sigma_{\delta d}}\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}_{\delta}(E)\right\}^{2}\right]<\epsilon^{2}.$ (D.18)
@@
$\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}_{s}(E)\right\}^{2}\right]<\lambda^{2},$ (D.19)
@@
$\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}_{\delta}(E)\right\}^{2}\right]\geq\lambda^{2}.$ (D.20)
Eq. (D.20) further implies
$\frac{1}{d}\sum_{\mathcal{H}_{\delta}(E)\subseteq\Sigma_{d}}\operatorname{Tr}\left[\left\{\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\hat{\Pi}_{\delta}(E)\right\}^{2}\right]\geq\frac{\lambda^{2}}{d}\left(\sum_{\mathcal{H}_{\delta}(E)\subseteq\Sigma_{\delta d}}1\right).$ (D.21)
Combining this with Eq. (D.18), we get
$\left(\sum_{\mathcal{H}_{\delta}(E)\subseteq\Sigma_{\delta d}}1\right)\leq\frac{\epsilon^{2}}{\lambda^{2}}d,$ (D.22)
which is Eq. (C.2).
@@
$\left\lvert\int\mathrm{d}t\ w(t)\operatorname{Tr}[\hat{\rho}(t)\hat{\Pi}_{A}]-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\right\rvert=\left\lvert\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}\rangle,\lvert E_{m}\rangle\in\Sigma_{d}\end{subarray}}\widetilde{w}(E_{n}-E_{m})\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\langle E_{m}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\lvert E_{n}\rangle\right\rvert$. (D.23)
To simplify our notation, we will take $\lvert E_{n}\rangle,\lvert E_{m}\rangle\in\Sigma_{d}$ as given in what follows, without explicitly writing out this condition in the sums. Separating Eq. (D.23) into contributions from within and outside the bandwidth $\Delta E$ (noting that $\hat{\mathds{1}}$ has no off-diagonal matrix elements) and using the triangle inequality, we get
[…]
$+\left\lvert\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\widetilde{w}(E_{n}-E_{m})\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\langle E_{m}\rvert\hat{\Pi}_{A}\lvert E_{n}\rangle\right\rvert$. (D.24)
@@
$\leq\sqrt{\left(\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\left\lvert\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\right\rvert^{2}\right)\operatorname{Tr}\left\{\left[\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right]^{2}_{(\Sigma_{d},\Delta E)}\right\}}$ (D.25)
@@
$\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\left\lvert\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\right\rvert^{2}\leq\operatorname{Tr}[\hat{\rho}^{2}].$ (D.26)
@@
$\left\lvert\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\widetilde{w}(E_{n}-E_{m})\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\langle E_{m}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\lvert E_{n}\rangle\right\rvert<\epsilon\sqrt{d\operatorname{Tr}[\hat{\rho}^{2}]}.$ (D.27)
@@
$\leq w_{0}\sqrt{\left\{\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\left\lvert\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\right\rvert^{2}\right\}\left\{\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\left\lvert\langle E_{m}\rvert\hat{\Pi}_{A}\lvert E_{n}\rangle\right\rvert^{2}\right\}}$ (D.28)
@@
$\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\left\lvert\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\right\rvert^{2}$ $\leq\operatorname{Tr}[\hat{\rho}^{2}]$ (D.29)
$\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\left\lvert\langle E_{m}\rvert\hat{\Pi}_{A}\lvert E_{n}\rangle\right\rvert^{2}$ $\leq\operatorname{Tr}\left[\left\{\hat{\Pi}_{A}\hat{\Pi}_{\Sigma_{d}}\right\}^{2}\right].$ (D.30)
@@
$\operatorname{Tr}\left[\left\{\hat{\Pi}_{A}\hat{\Pi}_{\Sigma_{d}}\right\}^{2}\right]\leq\operatorname{Tr}[\hat{\Pi}_{A}\hat{\Pi}_{\Sigma_{d}}]=d\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}},$ (D.31)
@@
$\left\lvert\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\widetilde{w}(E_{n}-E_{m})\langle E_{n}\rvert\hat{\rho}\lvert E_{m}\rangle\langle E_{m}\rvert\hat{\Pi}_{A}\lvert E_{n}\rangle\right\rvert\leq w_{0}\sqrt{d\operatorname{Tr}[\hat{\rho}^{2}]\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}}.$ (D.32)
Inserting Eqs. (D.27) and (D.32) into Eq. (D.24), we get Eq. (66).
@@
Noting that $\mathbbR\lambda_{k}[w]\in\mathbb{R}$ [defined in Eq. (74)], it is convenient to separate the basis states in $\mathcal{B}$ into a set of states $\mathcal{B}_{+}$ in which $\lambda_{k}[w]\geq 0$, numbering $n_{+}$, and a set $\mathcal{B}_{-}$ in which $\lambda_{k}[w]<0$, with $n_{-}$ states:
$\mathcal{B}_{+}$ $=\{\lvert k\rangle\in\mathcal{B}:\ \lambda_{k}[w]\geq 0\},\;\;\ \lvert\mathcal{B}_{+}\rvert=n_{+},$ (D.33)
$\mathcal{B}_{-}$ $=\{\lvert k\rangle\in\mathcal{B}:\ \lambda_{k}[w]<0\},\;\;\ \lvert\mathcal{B}_{+}\rvert=n_{-}.$ (D.34)
@@
$\hat{\rho}_{+}$ $\equiv\frac{1}{n_{+}}\sum_{\lvert k\rangle\in\mathcal{B}_{+}}\lvert k\rangle\langle k\rvert$, (D.35)
$\hat{\rho}_{-}$ $\equiv\frac{1}{n_{-}}\sum_{\lvert k\rangle\in\mathcal{B}_{-}}\lvert k\rangle\langle k\rvert$, (D.36)
@@
$\frac{1}{n_{\pm}}\sum_{k\in\mathcal{B}_{\pm}}\left\lvert\lambda_{k}[w]\right\rvert<\left(\epsilon+w_{0}\sqrt{\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}}\right)\sqrt{\frac{d}{n_{\pm}}},$ (D.37)
@@
$\frac{1}{d}\sum_{k\in\mathcal{B}_{\pm}}\left\lvert\lambda_{k}[w]\right\rvert<\left(\epsilon+w_{0}\sqrt{\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}}\right)\frac{\sqrt{n_{+}}+\sqrt{n_{-}}}{\sqrt{d}}\leq\left(\epsilon+w_{0}\sqrt{\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}}\right)\sqrt{2},$ (D.38)
@@
$f\lambda\leq\frac{1}{d}\sum_{k\in\mathcal{B}_{\pm}}\left\lvert\lambda_{k}[w]\right\rvert$, (D.39)
which, on substitution into Eq. (D.38), gives Eq. (76).
@@ D.4 Global thermalization from a single initial state
$+\frac{1}{\operatorname{Tr}[\hat{\Pi}_{A}]}\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert\geq\Delta E\end{subarray}}\widetilde{w}_{+}(E_{n}-E_{m})\left\lvert\langle E_{n}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle\hat{\mathds{1}}\right)\lvert E_{m}\rangle\right\rvert^{2}$ (D.40)
@@
$\int\mathrm{d}t\ w_{+}(t)\operatorname{Tr}[\hat{\rho}_{A}(t)\hat{\Pi}_{A}]-\langle\hat{\Pi}_{A}\rangle\geq\frac{1}{\operatorname{Tr}[\hat{\Pi}_{A}]}\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\widetilde{w}_{+}(E_{n}-E_{m})\left\lvert\langle E_{n}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle\hat{\mathds{1}}\right)\lvert E_{m}\rangle\right\rvert^{2}$ (D.41)
@@
$\frac{1}{D}\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\left\lvert\langle E_{n}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle\hat{\mathds{1}}\right)\lvert E_{m}\rangle\right\rvert^{2}<\frac{\operatorname{Tr}[\hat{\Pi}_{A}]}{WD}\left[\int\mathrm{d}t\ w_{+}(t)\operatorname{Tr}[\hat{\rho}_{A}(t)\hat{\Pi}_{A}]-\langle\hat{\Pi}_{A}\rangle\right].$ (D.42)